OpenStudy (anonymous):

Explain how to simplify 3/1 + 2/a^2 - 3/a^3 please?

OpenStudy (anonymous):

is this your qn. \[\frac{ 3 }{ 1}+ \frac{ 2 }{ a ^{2} } -\frac{ 3 }{ a ^{3}}\]

OpenStudy (anonymous):

Yes, it is.

OpenStudy (anonymous):

put everything over the same denominator (a^3) and then just add the numerator. It will end up being (3a^3 + 2a -3)/a^3